The combined cost of one advance ticket to a show and one same-day ticket was $55 . It is known that 30 tickets were sold in advance and 25 the same day, for total receipts of $1575 . What was the price of each kind of ticket?

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

30X+25(55-X)=1575 OR 30X+1375-25X=1575 OR 5X=200 OR X=200/5 OR X=40 TICKETS SOLD FOR $30 EACH & 55-40=15 TICKETS SOLD FOR $25
PROOF 30*40+15*25=1575 OR 1200+375=1575 OR 1575=1575

Answer by stanbon(75887)

The combined cost of one advance ticket to a show and one same-day ticket was $55 . It is known that 30 tickets were sold in advance and 25 the same day, for total receipts of $1575 . What was the price of each kind of ticket?
Let cost of advance tickets be "x"
Then cost of same day tickets is "55-x"
Value of advance tickets sold is 30x
Value of same day tickets sold is 25(55-x)
EQUATION:
Value + Value = $1575
30x+25(55-x)=1575
30x+(25)(55)-25x=1575
5x=1575-25(55)
x=315-275
x=$40 (price of advance tickets)
55-x=$15 (price of same day tickets)
Cheers,
Stan H.